A ship is traveling at 32 knots for 6 hours on a course of 155 degrees. How far south adn east did he travel?
Answer by Alan3354(69443)

A ship is traveling at 32 knots for 6 hours on a course of 155 degrees. How far south adn east did he travel?
------------------
155 = 65 degrees from the x-axis
distance = 32*6 = 192 mn
x = 192*cos(65) East
y = 192*sin(65) South